OpenStudy (anonymous):

What is the area of a circle with radius 14 units, to the nearest hundredth?

OpenStudy (anonymous):

\[\huge~\rm~A=\pi~r^2\] \[\huge~\rm~A=(3.14)~(14)^2\]

OpenStudy (anonymous):

615.44

OpenStudy (anonymous):

14^2=?

OpenStudy (anonymous):

Thats 196